OpenStudy (anonymous):

If sine of x equals 1 over 2, what is cos(x) and tan(x)? Explain your steps in complete sentences.

OpenStudy (anonymous):

\[sinx = \frac{ 1 }{ 2 }\] as such correct?

OpenStudy (anonymous):

@iambatman yes

OpenStudy (anonymous):

Notice the ratio for sinx is \[\sin(x) = \frac{ \text{opposite} }{ \text{hypotenuse} }\] so we can make a right triangle  we can use pythagorean theorem to find the adjacent side, can you do that please?

OpenStudy (anonymous):